OpenStudy (anonymous):

2. Planning is crucial in the early stages of this project. Demonstrate how a recursive process will allow you to find the number of coins and points on all levels up to level 5. The 3. The development team has asked you to jump ahead of them in the project. Create the sequence formulas, an, for the coins and the points based on the level in the game. Then describe how the formula can be used to find the coins and values on level 15. Use complete sentences.

OpenStudy (anonymous):

4. If the game only has 20 levels, explain how to find the value of the series for the coins and the points. Use complete sentences and arrive at final values. 5. The programming team needs to understand the parameters for the game, in order to start coding it. Explain any restrictions to the domain and range of your sequences.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

the answers to the first two were:1. Create the data to fill in the tables below. The Coins table must be an arithmetic sequence and the Points table must be a geometric sequence. The common difference or ratio cannot equal 1 or 0. 2,4,6,8 for coins. And 2, 4, 8, 16 for points. The common ratio is two. Provide reasons and justification of how you know the Coins sequence you created is arithmetic and how you know the Points sequence is geometric. I know that the coins in the sequence are arithmetic because I can add two each time. The points are geometric because I multiplied by two each time

OpenStudy (sleepyjess):

do you want me to put what i did up here and what my teacher said so you can see what to do?

OpenStudy (anonymous):

that would be great!!

OpenStudy (sleepyjess):

here is what i did

OpenStudy (sleepyjess):

here is what my teacher said about it Good work Jesstina, but #4-5 are incorrect. For #4, please be sure to uses the series formulas to find the sum and show work. For #5, the domain does not include all numbers, rather, it only includes integers.

OpenStudy (sleepyjess):

does that help?

OpenStudy (anonymous):

kind of, its weird because your numbers are different

OpenStudy (sleepyjess):

what lesson is the material in? i will try to help you with your numbers

OpenStudy (anonymous):

its in lesson 8 im really confused on the recursive process part, i don't even know what to do for that

OpenStudy (sleepyjess):

this is the recursive process for the coins A recursive process requires the use of the previous term in a sequence. "Add two to get the next number" is what you might say when thinking recursively. A recursive process could be listed as an = an – 1 + 2. The subscript n and n – 1 refer to the place in the sequence a term is. The first term is a1, the second term is a2, and so on. Using the recursive process an = an – 1 + 2 we can find the terms of an arithmetic sequence given the first term. Assume the first term, a1, equals 3.

OpenStudy (sleepyjess):

the recursive process formula would be \[a _{n}=a _{n-1}+a _{1}\]

OpenStudy (sleepyjess):

so for your numbers it would be a4 = a3 +2

OpenStudy (sleepyjess):

and a3 is 6

OpenStudy (anonymous):

so id just solve it down?

OpenStudy (sleepyjess):

yes then you would do the same thing for a5 but using a4 in place of a3 a5=a4+2 a4 being 8

OpenStudy (anonymous):

that's it?

OpenStudy (sleepyjess):

this is for the points A recursive process requires the use of the previous term in a sequence. “Multiply by four to get the next number” is what you might say when thinking recursively. A recursive process could be listed as an = 4(an – 1). The subscript n and n – 1 refer to the place in the sequence a term is. The first term is a1, the second term is a2, and so on. Using the recursive process an = 4(an – 1) we can find the terms of a geometric sequence given the first term. Assume the first term, a1, equals 3.

OpenStudy (sleepyjess):

but you are multiplying by 2 so it would be \[a _{n}=2(a _{n-1})\]

OpenStudy (sleepyjess):

are you still here?

OpenStudy (anonymous):

yeah

OpenStudy (sleepyjess):

do you get it so far?

OpenStudy (anonymous):

im trying to .it makes enough sense, so after i get the points i use that^

OpenStudy (sleepyjess):

that is to get the points for level 4 your equation would be \[a _{4}=2(a _{3})\] a3=8

OpenStudy (sleepyjess):

for level 5 you would just replace a4 with a5 and a3 with a4 which would be....?

OpenStudy (anonymous):

i don't know. im confused. with what im supposed to be putting in

OpenStudy (sleepyjess):

do you get the coins formula?

OpenStudy (anonymous):

i honestly don't get any of it. the most i really understand is how to get the numbers i had.. im going back through the lessons now to see if i can make sense of it